I need help on part C
Problem 7.69 Constants | Periodic Table elevator cable breaks when a 975 kg elevator is 24.5 m above the top of a huge spring (k 6.50-104 N/m) at the bottom of the shaft
Part A Calculate the work done by gravity on the elevator before it hits the spring WG 2.34x105 J Previous Answers Correct Part B Calculate the speed of the elevator just before striking the spring. v 21.9 m/s Previous Answers Correct
Part C Calculate the amount the spring compresses (note that here work is done by both the spring and gravity) 2.67 Submit Previous Answers Reauest Answer X Incorrect; Try Again; 3 attempts remaining
Answer #1

work done by gravity = 975*9.81* 24.5= 234336.4

The work stops when work done by gravity equals the work done to stop it by the spring (Kinetic energy).

Kinetic energy= Elastic energy stored by spring = 1/2*k*(compression^2)

= 1/2 *6.50x104 *x2

1/2 *6.50x104 *x2 = 234336.4

3.25x104x2 = 234336.4

x2 = 7.210

x=1.6386m